### v3

@@ II A brief revision of Lovelock gravity with $n$ fold degenerated ground state
where $R^{\mu\alpha}_{\hskip 6.02777pt\nu\beta}$ is the Riemann tensor, $[X]$ stands for the integer part of $X$ and $\{\alpha_{p}\}$ is a set of dimensionful constants. The LL equations of motion can be written as
@@
Unfortunately, not even in (matter) vacuum these equations of motion can be solved straightforwardly for an arbitrary set of $\alpha$’s. Not even the ground-states, i.e. the locally constant curvature solutions, can be obtained in $d>8$ for a general set of $\alpha$’s. The problem of obtaining those constant curvatures solutions reduces to obtain the roots of a polynomial of order $[d/2]$, which, as it is well known, is not unattainable for $[d/2]>4$. Moreover, the roots can be complex numbers with nonvanishing imaginary parts. Therefore, to remove any problem, the asymptotic behavior is fixed to a single asymptotically locally flat or AdS behavior.
@@
[…]
@@
[…]
@@
$ds^{2}=-f(r)^{2}dt^{2}+\frac{1}{f(r)^{2}}dr^{2}+r^{2}d\Sigma_{\gamma}$ (7)
where $\Sigma_{\gamma}$ is a manifold of constant curvature $\gamma=\pm 1,0$, and where the vacuum solution is
$f(r)^{2}=\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}.$ (8)
It must be stressed that $(X)^{1/n}$ stands for any of the $n$ different roots of (X). It is straightforward to check the presence of a singularity at $r=0$. Nonetheless, this is not enough, and in general, what characterizes the interesting physical solutions is that they satisfy $f(r)^{2}\in\mathbb{R}$, for $r>0$. If ∃$\exists\hskip 4.30554ptr^{*}$ such $f(r^{*})^{2}=0$ this defines a Killing horizon and thus the line element above Eqs.(7) and (8) describe a black hole geometry. One can notice how similar is this solution to the corresponding Schwarzschild solution  [^1] Nonetheless, at least there is significant difference. For $n$ even integer, since there are always two $\pm$ real roots in the expression, $M<0$ can be considered allowed..
It is worth stressing that, as mentioned in the introduction, the structure of the equations of motion leading to the last solution is quite different from that of PL, as PL considers only a single term in the Lovelock series, while LnFDGS includes all the terms of the mentioned series up to $n=[d/2]$ under the particular choice of coupling constants previously mentioned. This choice allows the equations of motion to be factorized into AdS vacua, as described above (which is not possible for PL).
III About the nature of the matter source proposed for Lovelock with $n$ fold degenerated ground state AdS.
Given the aforementioned differences between PL and LnFDGS theories, providing a form for the energy density analogous to the one described earlier is not straightforward. This is because several additional terms would appear when directly substituting the Kretschmann scalar of the vacuum LnFDGS solution. To address this, we will provide some ingredients. First, we will define gravitational tension through the gauge curvature tensor $F_{\mu\nu}$, which is the Poincaré (group) curvature. Second, we will define an appropriate tensor for LnFDGS, which we will call the Poincaré AdS-like curvature. From this, we will provide an alternative version of the Kretschmann scalar associated with gravitational tension. Finally, we will describe, from a physical perspective, the resulting structure of the energy density.
@@
Although the analogy between gravitational tension and the production ratio in the Schwinger effect requires further study, in this work we can speculate by providing some glimpses on the subject. As discussed in Ansoldi (2008); Dymnikova (1992), one could argue that the Schwinger effect should play a role in one of two ways. Firstly, it would give rise to corrections to the vacuum expectation value of the fields. These corrections can be expressed in terms of powers of the Riemann tensor Birrell and Davies (1982). Furthermore, gravitation itself should self-correct due to the presence of singularities. It is expected that these effects cannot be separated for large curvature. To address any of these approximations, after considering Eq. (1), one can recall that in asymptotically (Riemannian) flat spaces (i.e., with a vanishing cosmological constant), one analogy to the gauge curvature $F_{\mu\nu}$ is the Poincaré (group) curvature. In this way, we can also interpret the gravitational tension as $F=\sqrt{F_{\mu\nu}F^{\mu\nu}}$, where
@@
To extend the framework to asymptotically (locally) Anti de Sitter spaces, it is enough to consider a $SO(d-1,2$) curvature Hassaine and Zanelli (2016) instead of the Poincare one. One option is Hassaine et al. (2005)
@@
In this work, for simplicity, we will refer to the tensor $F^{\text{AdS}}_{\mu\nu}$ as Poincaré AdS-like curvature, where $l$ is called the AdS radius and $J_{ab},J_{a}$ are a set of generators of $SO(d-1,2)$. Since the discussion is for Riemannian manifold, where $T^{\rho}_{\hskip 3.01389pt\mu\nu}=0$, in order to simplify the calculations, one can define an adequate Kretschmann scalar $K^{\prime}$
@@ III.3 The proposed model
In this way, the steps previously taken, namely: defining a tensor analogous to the gauge curvature, which we have called as PoincarĂŠ AdS-like curvature and providing an alternative definition for the Kretschmann scalar, now allow us to model an energy density. Thus, in analogy with (1), it can be defined the energy density
@@ IV A new static regular solution in Lovelock with $n$ fold degenerated ground state AdS
$\frac{d}{dr}\left(r^{d-1}\left(\bar{f}_{\gamma}(r)+\frac{1}{l^{2}}\right)^{n}\right)=\frac{2}{d-2}r^{d-2}\rho,$ (22)
@@
$\bar{f}_{\gamma}(r)=\frac{\gamma-f(r)^{2}}{r^{2}}$ (23)
On the other hand, due to the transverse symmetry, we have $p_{\theta}=p_{\phi}=p_{t}=\ldots$ for all the $(d-2)$ angular coordinates. Thus, using the aforementioned condition $\rho=-p_{r}$, the conservation law $T^{\alpha\beta}_{;\beta}=0$ gives:
@@
$f(r)^{2}=\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2m(r)}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (25)
@@
[…]
@@
$=$ $\frac{M}{n}{}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)\left(\frac{r}{a}\right)^{d-1}$
$=$ $M(n-1)!\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)$
where ${}_{1}F_{1}$ is hyperbolic confluent function. See appendix A.
Finally, the solution can be written as
$f(r)^{2}$ $=$ $\gamma+\frac{r^{2}}{l^{2}}-\frac{r^{2}}{a^{2}}\left(\frac{2M}{n}{}_{1}F_{1}\left(n,n+1,-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)\right)^{\frac{1}{n}}$ (28)
$=$ $\gamma+\frac{r^{2}}{l^{2}}-\left(\frac{2M(n-1)!}{r^{d-2n-1}}\left(1-\exp{\left(-\left(\frac{r}{a}\right)^{\frac{d-1}{n}}\right)}\sum^{n-1}_{i=0}\frac{1}{i!}\left(\frac{r}{a}\right)^{\frac{i}{n}(d-1)}\right)\right)^{\frac{1}{n}}$
Before proceeding it is worthwhile noticing the coordinate $r$ defines the regions of interest. While $r\rightarrow\infty$ defines the asymptotic region, $r\rightarrow 0$ must define the center of the solution. Second, this solution has two parameters of distance competing, $l$ and $a$, and thus one can foresee that $a/l$ should have a relevant rôle in the analysis. Furthermore, given respective physical meaning of $a$ and $l$, in a physically sound scenario it must be satisfied $a/l\ll 1$. Because of that one can expect that if an event horizon exist, then the value of the radius of the external horizon, says $r_{+}$, should have only small deviation from one associated to the vacuum solution found in Aros et al. (2001).
One central feature of (28) is that any analysis of its behaviour can only be done numerically due to the hypergeometric function. This is quite forlon as in this case the numerical analysis bypasses the small corrections due to $a/l\ll 1$, the sound physical scenario. On the other hand, the silver lighting is that model simultaneously solves the singularities while proposing black hole geometries that are, in practice, indistinguishable from their (singular) counterparts.
IV.1 Behavior near $r=0$ region
@@
Firstly one has to analyze the region near $r=0$, which basically separates the regular solution from its black hole counterpart. At $r=0$, one can show the lack of any singularity. This can be noticed by merely taking the limit
$\lim_{r\rightarrow 0}f(r)^{2}\rightarrow\gamma-\left(\left(\frac{2M}{n}\right)^{\frac{1}{n}}-\frac{a^{2}}{l^{2}}\right)\frac{r^{2}}{a^{2}}.$ (29)
In principle this defines either a (nearly) (Anti) de Sitter or a flat region nearby $r=0$. However, one can notice the presence of $a^{2}/l^{2}$ and thus, because in a realistic physical set-up $a/l\ll 1$, even for small masses $M$, recalled measured in term of units of $a^{d-3}$, it should be expected a de Sitter space region near $r=0$. This recovers the behavior without cosmological constant. Regarding the latter, in the following of this work, we will consider $\gamma=1$, i.e., the case where the transversal section is spherically symmetric.

### v4

@@ II A brief revision of Lovelock gravity with n-fold degenerated ground state
where $R^{\mu\alpha}_{\hskip 6.02777pt\nu\beta}$ is the Riemann tensor, $[X]$ stands for the integer part of $X$ and $\{\alpha_{p}\}$ is a set of dimensionful constants. As mentioned in the introduction, since they are dimensionful coupling constants, they can be defined such that they have units of $\alpha_{p}\sim[\text{length}]^{2p-d}$ Garraffo and Giribet (2008). Thus, it is possible to define $\alpha_{0}\sim\Lambda\alpha_{n}$, with $\Lambda$ having units of $\Lambda\sim[\text{length}]^{-2n}$. Therefore, for $n=1$, corresponding to the Einstein-Hilbert case, we have $\Lambda\sim[\text{length}]^{-2}$. The LL equations of motion can be written as
@@
It’s straightforward to notice that not even in a vacuum can these equations of motion be solved for an arbitrary set of $\alpha$’s. Not even the ground-states, i.e. the locally constant curvature solutions, can be obtained in $d>8$ for a general set of $\alpha$’s. The problem of obtaining those constant curvature solutions reduces to getting the roots of a polynomial of order $[d/2]$, which, as it is well known, is not unattainable for $[d/2]>4$. Moreover, those roots can be complex numbers with nonvanishing imaginary parts. A way to avoid that discussion is to fix the asymptotic behavior to a single asymptotically locally flat or AdS behavior.
@@
$\mathcal{G}^{\mu}_{(LL)\hskip 3.01389pt\nu}\propto\delta_{\mu_{1}\nu_{1}\ldots\mu_{I}\nu_{I}\nu}^{\alpha_{1}\beta_{1}\ldots\alpha_{I}\beta_{I}\mu}(R^{\nu_{1}\mu_{1}}_{\hskip 6.02777pt\alpha_{1}\beta_{1}}+\kappa_{1}\delta^{\mu_{1}\nu_{1}}_{\hskip 6.02777pt\alpha_{1}\beta_{1}})\ldots(R^{\nu_{I}\mu_{I}}_{\hskip 6.02777pt\alpha_{I}\beta_{I}}+\kappa_{I}\delta^{\mu_{I}\nu_{I}}_{\hskip 6.02777pt\alpha_{I}\alpha_{I}})=0.$ (4)
@@
$\mathcal{G}^{\alpha}_{\hskip 3.01389pt\beta}=\alpha_{0}\delta^{\alpha\nu_{1}\ldots\nu_{2n}}_{\beta\mu_{1}\ldots\mu_{2n}}\left(R^{\mu_{1}\mu_{2}}_{\hskip 6.02777pt\nu_{1}\nu_{2}}+\frac{1}{l^{2}}\delta^{\mu_{1}\mu_{2}}_{\nu_{1}\nu_{2}}\right)\ldots\left(R^{\mu_{2p-1}\mu_{2p}}_{\hskip 6.02777pt\nu_{2n-1}\nu_{2n}}+\frac{1}{l^{2}}\delta^{\mu_{2n-1}\mu_{2n}}_{\nu_{2n-1}\nu_{2n}}\right)$ (6)
@@
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}d\Omega_{d-2}$ (7)
where $d\Omega_{d-2}$ corresponds to the transversal section of a (d2)(d-2)-sphere. The vacuum solution is
$f(r)=1+\frac{r^{2}}{l^{2}}-\left(\frac{2M}{r^{d-2n-1}}\right)^{\frac{1}{n}}.$ (8)
It must be stressed that $(X)^{1/n}$ stands for any of the $n$ different roots of (X). It is straightforward to check the presence of a singularity at $r=0$. Nonetheless, this is not enough, and in general, what characterizes the interesting physical solutions is that they satisfy $f(r)^{2}\in\mathbb{R}$, for $r>0$. If $\exists\hskip 4.30554ptr^{*}$ such $f(r^{*})^{2}=0$ this defines a Killing horizon and thus the line element above Eqs.(7) and (8) describe a black hole geometry. One can notice how similar is this solution to the corresponding Schwarzschild solution  [^1] Nonetheless, at least there is significant difference. For $n$ even integer, since there are always two $\pm$ real roots in the expression, $M<0$ can be considered allowed..
It is worth stressing that, as mentioned in the introduction, the structure of the equations of motion leading to the last solution is quite different from that of PL, as PL considers only a single term in the Lovelock series. At the same time, LnFDGS includes all the terms of the mentioned series up to $n=[d/2]$ under the particular choice of coupling constants previously mentioned. This choice allows the equations of motion to be factorized into AdS vacua, as described above (which is not possible for PL).
III About the nature of the matter source proposed for Lovelock with N-fold degenerated ground state AdS.
Given the aforementioned differences between PL and LnFDGS theories, providing a form for the energy density analogous to the one described earlier is not straightforward. This is because several additional terms would appear when directly substituting the Kretschmann scalar of the vacuum LnFDGS solution. To address this, we will provide some of the necessary ingredients. First, we will define gravitational tension through the gauge curvature tensor $F_{\mu\nu}$, which is the Poincaré (group) curvature. Second, we will define an appropriate tensor for LnFDGS, which we will refer to as the Poincaré AdS-like curvature. From this, we will provide an alternative version of the Kretschmann scalar associated with gravitational tension. Finally, we will describe, from a physical perspective, the resulting structure of the energy density.
@@
Although the analogy between gravitational tension and the production ratio in the Schwinger effect requires further study, this work provides some insights into the subject, allowing us to speculate. As discussed in Ansoldi (2008); Dymnikova (1992), one could argue that the Schwinger effect should play a role in one of two ways. Firstly, it would give rise to corrections to the vacuum expectation value of the fields. These corrections can be expressed in terms of powers of the Riemann tensor Birrell and Davies (1982). Furthermore, gravitation itself should self-correct due to the presence of singularities. It is expected that these effects cannot be separated for large curvature. To address any of these approximations, after considering Eq. (1), one can recall that in asymptotically (Riemannian) flat spaces (i.e., with a vanishing cosmological constant), one analogy to the gauge curvature $F_{\mu\nu}$ is the Poincaré (group) curvature. In this way, we can also interpret the gravitational tension as $F=\sqrt{F_{\mu\nu}F^{\mu\nu}}$, where
@@
To extend the framework to asymptotically (locally) Anti de Sitter spaces, it is enough to consider a SO(d1,2SO(d-1,2) curvature Hassaine and Zanelli (2016) instead of the Poincare one. One option is Hassaine et al. (2005)
@@
In this work, for simplicity, we will refer to the tensor $F^{\text{AdS}}_{\mu\nu}$ as Poincaré AdS-like curvature, where $l$ is called the AdS radius and $J_{ab},J_{a}$ are a set of generators of SO(d1,2)SO(d-1,2). Since the discussion is for Riemannian manifold, where $T^{\rho}_{\hskip 3.01389pt\mu\nu}=0$, in order to simplify the calculations, one can define an adequate Kretschmann scalar $K^{\prime}$
@@ III.3 The proposed model
In this way, the steps previously taken, namely: defining a tensor analogous to the gauge curvature, which we have called as Poincaré AdS-like curvature and providing an alternative definition for the Kretschmann scalar, now allow us to model an energy density. Thus, in analogy with (1), it can be defined the energy density
@@ IV A new static regular solution in Lovelock with $n$ fold degenerated ground state AdS
$\frac{d}{dr}\left(r^{d-1}\left(\bar{f}(r)+\frac{1}{l^{2}}\right)^{n}\right)=\frac{2}{d-2}r^{d-2}\rho,$ (22)
@@
$\bar{f}(r)=\frac{1-f(r)}{r^{2}}$ (23)
On the other hand, due to the transverse symmetry, we have $p_{\theta}=p_{\phi}=p_{t}=\ldots$ for all the (d2)(d-2) angular coordinates. Thus, using the aforementioned condition $\rho=-p_{r}$, the conservation law $T^{\alpha\beta}_{;\beta}=0$ gives:
@@
$f(r)=1+\frac{r^{2}}{l^{2}}-\left(\frac{2m(r)}{r^{d-2n-1}}\right)^{\frac{1}{n}}$ (25)
@@
$m(r)=\frac{1}{d-2}\int\rho r^{d-2}dr$ (26)
We replace the energy density in the last equation, using equations (18) and (19), obtaining the mass function
@@
In the expression above, we have chosen $M\cdot(n-1)!$ as the integration constant. As we will see below, this choice allows the function to behave as a de Sitter core close to the origin, and thus, the solution is regular.
It is worth noting that the coordinate $r$ defines the regions of interest. While $r\rightarrow\infty$ corresponds to the asymptotic region, $r\rightarrow 0$ must define the center of the solution. Below, we will analyze the regions associated with the event horizons present in the solution.
IV.1 Structure of horizons
We will begin by mentioning that the mass parameter, at which our solution, equations (25) and (26) vanishes, is given by the following expression:
$M(\bar{r})=\frac{\bar{r}^{-1+d-2n}\left(1+\frac{\bar{r}^{2}}{l^{2}}\right)^{n}}{2\left((-1+n)!-\Gamma\left[n,\left(\frac{\bar{r}}{a}\right)^{\frac{d-1}{n}}\right]\right)}$ (28)
where $\bar{r}$ is such that $f(r=\bar{r})=0$.
In Figure 1, we can observe the generic behavior of the mass parameter as a function of the radial values where the function $f(r)$ vanishes. The behavior in the vacuum LnFDGS AdS case Crisostomo et al. (2000); Aros et al. (2001) is shown in dashed red, while the behavior of our RBH LUV AdS solution with our matter sources is shown in blue. Thus, the regions where $M$ decreases correspond to the values of the inner horizon $\bar{r}=r_{-}$, and the regions where $M$ increases correspond to the values of the black hole horizon $\bar{r}=r_{+}$ in our solution. Therefore, for $M>M_{\text{min}}$, there exist an inner horizon $r_{-}$ and a black hole horizon $r_{+}$ for the same value of the parameter $M$. The value $M_{\text{min}}$ corresponds to the point where the inner and black hole horizons coincide, i.e., an extremal black hole.
A remarkable feature is that, under a small rightward deviation in the value of the extremal radius $r_{\text{ext}}=r_{-}=r_{+}$, that is, $\bar{r}=r_{*}>r_{\text{ext}}$ (which corresponds to a small upper-right deviation in the value of $M$ with respect to $M_{\text{min}}$), our regular AdS solution and its vacuum AdS counterpart become indistinguishable, as the dashed red and blue curves also become indistinguishable. More specifically, for a value slightly above $M_{\text{min}}$, both the vacuum solution and the regular solution share the same event horizon radius. In this case, the two solutions can only be distinguished inside the event horizon.
Figure 1: In horizontal axis $\bar{r}$. In vertical axis: $M$ of our RBH is shown in blue and $M$ in the vacuum case Aros et al. (2001) in dashed red. We have used $d=6,n=2,l=10,a=1$. We can check that this behavior is generic by using other parameter values. The decreasing part of the blue curve corresponds to the inner horizon of the RBH solution, while the increasing part represents the event horizon. The minimum point marks the extremal radius. The red curve represents the vacuum AdS solution. For a value slightly greater than the extremal radius, both solutions become indistinguishable for radial values equal to or greater than the event horizon.
For values on the vertical axis smaller than $M_{\text{min}}$, our solution exhibits a regular geometry without horizons. It is worthwhile to stress that, unlike the vacuum LnFDGS AdS solutions in Crisostomo et al. (2000); Aros et al. (2001), due to the absence of singularities, regular solutions do not need to be black holes, i.e., they do not require the presence of a horizon to be well defined. In other words, regular solutions are not constrained by the cosmic censorship hypothesis.
If we take the Planck length as a reference on the horizontal axis of the figure, such that the extremal radius is $r_{\text{ext}}\sim 1.6,\ell_{p}$ and $M_{\text{min}}\sim\ell_{p}$, then for values slightly greater than these, $r_{*}>r_{\text{ext}}$, both solutions become indistinguishable. In this scenario, physics at Planck scales becomes particularly interesting, as the nature of the matter sources contributes to the suppression of the singularity.
As a consequence of the above, we also note that for values considerably larger than the Planck scales, such as the observed values of mass and radius of rotating black holes observed in LIGO GW150914, about 35 solar masses Abbott et al. (2017b), both the vacuum and regular solutions are also indistinguishable. Thus, it becomes intriguing to observe the properties of black holes at scales much smaller than those that have been observed to date.
It is worth discussing the potential case where $r_{+}\gg a$, in which, as a consequence of the previous discussion, the Schwarzschild LnFDGS AdS solution Crisostomo et al. (2000); Aros et al. (2001) becomes indistinguishable from our regular black hole (RBH) solution with LnFDGS AdS. This scenario makes sense if one considers the case in which the parameter $a$ is of the order of the Planck length, $a\sim\ell_{p}$, and the radius is of the order of the rotating black hole observed in LIGO GW150914 Abbott et al. (2017b), $r_{+}\sim 103$ km. In this way, the parameter $a$ can be interpreted essentially as the length scale at which quantum corrections become not only non-negligible but also relevant. Depending on the model considered, this length scale can be as small as the Planck length.
IV.2 Merging zone
A remarkable feature of our regular solution is the existence, in general, of a radial coordinate value at which the regular solution converges toward the vacuum LnFDGS AdS solution found in Refs.Crisostomo et al. (2000); Aros et al. (2001), which corresponds to Eq.(8). However, analytically estimating the precise radial value beyond which our regular AdS solution and its vacuum AdS counterpart become indistinguishable is difficult. Numerically, we observe in the first panel of Figure 2 that, for a value of the mass parameter $M_{min}$, associated with the existence of an extremal horizon $r_{ext}$, both solutions become indistinguishable for a radial value greater than the extremal horizon. In the second panel, we observe that for mass values greater than $M_{min}$, corresponding to an event horizon $r_{*}=r_{+}>r_{ext}$, both solutions become indistinguishable before reaching the mentioned event horizon. It is straightforward to verify that the behavior shown in this figure is generic for other parameter choices. In this way, the obtained solution is such that, for a mass parameter value slightly greater than $M_{\text{min}}$ (associated with an extremal radius), our RBH solution LnFDGS becomes indistinguishable from its vacuum counterpart LnFDGS Crisostomo et al. (2000); Aros et al. (2001) inside the event horizon, from a radial value $r=r_{*}<r_{+}$ up to infinity. If $r_{*}$ is of the order of the Planck scale, the geometric differences between both solutions, particularly the suppression of the singularity, would occur at quantum scales.
Figure 2: On the horizontal axis, $r$. On the vertical axis, the function $f(r)$ for $M=M_{\text{ext}}\approx 1.01$, $M=1.2>M_{\text{ext}}$, and $M=2>M_{\text{ext}}$ in the first, second, and third panels, respectively. We have used $d=6$, $n=2$, $l=10$, $a=1$. We can check that this behavior is generic for other parameter values. We observe that for mass parameter values greater than the extremal mass, $M>M_{\text{ext}}$, both solutions become indistinguishable from within the event horizon.
IV.3 Behavior near $r=0$ region
@@
One must analyze the region near $r=0$, which essentially separates the regular solution from its black hole counterpart. At $r=0$, the absence of any singularity can be demonstrated. This becomes evident simply by taking the limit.
$\lim_{r\rightarrow 0}f(r)\rightarrow 1-\left(\left(\frac{2M}{n}\right)^{\frac{1}{n}}-\frac{a^{2}}{l^{2}}\right)\frac{r^{2}}{a^{2}}.$ (29)
n principle, this defines a (nearly) (Anti) de Sitter or flat region near $r=0$. However, the presence of the term $a^{2}/l^{2}$ can be observed, and therefore, given that in a realistic physical setup $a/l\ll 1$, where $l$ represents the AdS radius, even for small masses $M$, a de Sitter-like region is expected near $r=0$.